OpenStudy (anonymous):

elimination method solve for x and y 8x-9y=8.5 7y-2x=-4.5

OpenStudy (anonymous):

multipy the second equation by 4 to get 28y -8x = -18. the 8x and the -8x will cancel out and all you have to do is solve the leftover equation.

jhonyy9 (jhonyy9):

8x-9y=8,5 for you can eliminat the term x you need multiplie the second -2x+7y=-4,5 equation with 4 and hence you will get -8x+28y =-18 than you have 8x-9y=8,5 -8x+28y=-18 assum that and you get ------------------- 0 19y= -9,5 y= -9,5/19 = -1/2 - than 8x-9(-1/2)=8,5 8x+(9/2)=8,5 8x=8,5-(9/2)=(17-9)/2 =8/2=4 x=4/8 =1/2 x=1/2 y=-1/2